OpenStudy (anonymous):

Add the following fraction : 2/9+15/12+(-13/8)

OpenStudy (anonymous):

\[\frac{ 2 }{ 9 }+ \frac{ 15 }{ 12 }+ (- \frac{ 13 }{ 8 })\]

OpenStudy (blurbendy):

2 / 9 + 5 / 12 - 13 / 8 Start by adding 2 / 9 and 15 / 12 (you'll need the least common denominator)

OpenStudy (anonymous):

The cd is 1 isnt it?

OpenStudy (anonymous):

@blurbendy

OpenStudy (anonymous):

I would just plug it into your calculator or multiply each to find a common denominator

OpenStudy (blurbendy):

It would not be 1. You're looking for the smallest number 9 and 12 have in common when you keep going through their multiples. so 9, 18, 27 ... 12, 24 ....

OpenStudy (blurbendy):

if you take one more step in each of those sequences youll find the lcm!

OpenStudy (anonymous):

o i thought you meant for all 3 # s 9*4=34 & 12*4=36 so ,36

OpenStudy (blurbendy):

yup, 36! So you multiplied 9 by 4 to get to 36, so you need to multiply 2 / 9 by 4 on the top and bottom

OpenStudy (blurbendy):

and multiply the top and bottom of 15/12 by 3 (since you multiplied 12 by 3 to get to 36)

OpenStudy (anonymous):

8/36+15/36+ (- 13/8)

OpenStudy (blurbendy):

close, you forgot to multiply 15 by 3

OpenStudy (anonymous):

Yea i meant 45 sorry

OpenStudy (blurbendy):

Excellent! So you have 8 / 36 + 45 / 36. Now that we have a common denominator, just add the top across and leave the 36 alone

OpenStudy (anonymous):

53/36

OpenStudy (blurbendy):

Nice. We have 53 / 36 - 13 / 8. Uh oh, we need to find another common denominator! (hint what is 36 + 36?)

OpenStudy (anonymous):

is the answer 181/72

OpenStudy (blurbendy):

nope, but 72 is the common denominator for 53 / 36 -13 / 8

OpenStudy (blurbendy):

We multiplied 36 by 2 to get to 72, so multiply the top and bottom of 53 / 36 by 2

OpenStudy (anonymous):

11/18?

OpenStudy (blurbendy):

Nope, 53 *2 = 106 so 106 / 72 - 117 / 72 <--- We multiplied the top and bottom of 13 / 8 by 9.

OpenStudy (blurbendy):

Your answer is 106 / 72 - 117 / 72 which is.... ?